OpenStudy (anonymous):

prove that (5^n )-4n +1 is divisble by 16 for n ={0, 1, 2, 3,...n, n+1}

OpenStudy (anonymous):

plz chec k the question.. it may be (5^n )-4n -1 if so express it as (4+1)^n -4n -1 and use binomial theorems then 4^2 will come as factor

OpenStudy (anonymous):

correction n={1, 2, .... n, n+1}

OpenStudy (anonymous):

still quest is not correct it may be (5^n )-4n -1

OpenStudy (anonymous):

I think it is 5^n-4n-1 and n={2,3,4,5...}

OpenStudy (anonymous):

if n starts from 1 that will also do

OpenStudy (anonymous):

ok, 0/16=0

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

we use method inductive

OpenStudy (anonymous):

induction can also be used

OpenStudy (anonymous):

n=0 it is divisble by 16 suppose n=k it true define 5^k-4k-1 divisble by 16 we prove n=k+1 true 5^(k+1)-4(k+1)-1=5*5^k-4k-4-1=5*(5^k-4k-1)+16k we have 5*(5^k-4k-1) divisble by 16 and 16k divisble by 16 thus 5*(5^k-4k-1)+16k divisble by 16

OpenStudy (anonymous):

yes the ? is (5^n) -4n-1 , i am confused using binomial thereom mainly, because i have never used it.

OpenStudy (anonymous):

use P.M.I. instead

OpenStudy (kinggeorge):

Just to make what chanh wrote a bit easier to read, we have \[\large 5^{n+1}-4(n+1)-1=5^{n+1}-4n-5\]\[=5^{n+1}-20n-5+16n\]\[=5(5^n-4n-1)+16n\]Both sides are divisible by 16, so we're done.

OpenStudy (anonymous):

I use induction

OpenStudy (anonymous):

so you are allowed to only multiply -4n by 5 and to add 16n,

OpenStudy (anonymous):

KingGeorge explained we have -20n+16n=-4n

OpenStudy (anonymous):

oh, i see